Direct Materials and Direct Labor Variance Analysis Abbeville Fixture Company manufactures faucets in a small manufacturing...

  1. Direct Materials and Direct Labor Variance Analysis

    Abbeville Fixture Company manufactures faucets in a small manufacturing facility. The faucets are made from brass. Manufacturing has 40 employees. Each employee presently provides 35 hours of labor per week. Information about a production week is as follows:

    Standard wage per hour $13.20
    Standard labor time per faucet 10 min.
    Standard number of lbs. of brass 1.2 lbs.
    Standard price per lb. of brass $11.75
    Actual price per lb. of brass $12.00
    Actual lbs. of brass used during the week 10,753 lbs.
    Number of faucets produced during the week 8,700
    Actual wage per hour $13.6
    Actual hours for the week (40 employees × 35 hours) 1,400 hrs.

    Required:

    a. Determine the A detailed estimate of what a product should cost.standard cost per unit for direct materials and direct labor. Round the cost per unit to two decimal places.

    Direct materials standard cost per unit $
    Direct labor standard cost per unit $
    Total standard cost per unit $

    b. Determine the direct materials Price variance is the difference between the actual and standard prices, multiplied by the actual quantity.price variance, direct materials The cost associated with the difference between the standard quantity and the actual quantity of direct materials used in producing a commodity.quantity variance, and total direct materials The difference between actual cost and standard cost at actual volumes.cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Materials Price Variance $ Unfavorable
    • Favorable
    • Unfavorable
    Direct Materials Quantity Variance $ Unfavorable
    • Favorable
    • Unfavorable
    Total Direct Materials Cost Variance $ Unfavorable
    • Favorable
    • Unfavorable

    c. Determine the direct labor The cost associated with the difference between the standard rate and the actual rate paid for direct labor used in producing a commodity.rate variance, direct labor The cost associated with the difference between standard and actual hours of direct labor spent for producing a commodity.time variance, and total direct labor cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Labor Rate Variance $ Unfavorable
    • Favorable
    • Unfavorable
    Direct Labor Time Variance $ Favorable
    • Favorable
    • Unfavorable
    Total Direct Labor Cost Variance $ Favorable
    • Favorable
    • Unfavorable

    Feedback

    Unfavorable variances can be thought of as increasing costs (a debit). Favorable variances can be thought of as decreasing costs (a credit).

    Learning Objective 2, Learning Objective 3.

Loading item

There was an error loading this item. If this continues to occur, please contact Technical Support.

Check My Work

  • Previous
  • Next
  • 100% Correct
  • Partially Correct
  • Incorrect
  • Needs Instructor Grading

Basic Calculatorclose

-7040

UseEntBSBSpCEHomCEnd

789+

456-

123*

0.=/
Answer #1
a
Direct materials standard cost per unit 14.10 =1.2*11.75
Direct labor standard cost per unit 2.20 =13.2*(10/60)
Total standard cost per unit 16.30
b
Direct Materials Price Variance 2688 Unfavorable =10753*(12-11.75)
Direct Materials Quantity Variance 3678 Unfavorable =11.75*(10753-8700*1.2)
Total Direct Materials Cost Variance 6366 Unfavorable =(10753*12)-(8700*1.2*11.75)
c
Direct Labor Rate Variance 560 Unfavorable =1400*(13.6-13.2)
Direct Labor Time Variance -660 Favorable =13.2*(1400-8700*10/60)
Total Direct Labor Cost Variance -100 Favorable =(1400*13.6)-(8700*13.2*10/60)